# Evolving a Multi-Population Evolutionary-QAOA on Distributed QPUs

### v2

@@ III Set-up
The simulations for the experiments were done on the Scafell Pike high-performance computing system, hosted by Sci-Tech Daresbury (UK), equipped with dual-socket nodes with a 16-core Intel Xeon Gold CPU per socket. The hardware runs were performed on IBM’s 127-qubit Eagle processors ibm_cusco and ibm_nazca, which have an Error Per Layered Gate (EPLG) for a 100-qubit chain [33] of $5.9\%$ and $3.2\%$ respectively. Within these processors, the QAOA was set to run with $10^{4}$ shots per circuit. For reproducibility, we summarize all the parameters used in Tab. I, and they remain the same for all experiments unless stated otherwise.
TABLE I: List of all variables used, their description and values assigned for the hybrid E-QAOA
@@
$\gamma$, $\beta$ Ansatz parameters to optimize $(-\pi$, $\pi]$
@@ IV Results and Discussion
Now, we present the results of our experiments and a discussion on their performance. Tests were first run on a simulator to quickly evaluate potential advantages and tune the appropriate parameters. Then, they were run on quantum hardware.
@@ IV-A Simulations
In Fig. 2 we show the performance of our EA optimizer against COBYLA, for increasing graph sizes. On the top plot, we note that for small graphs both methods achieve the optimal solution, i.e. an approximation ratio of 1. Whereas from $n\geq 16$, both optimizers start decreasing in accuracy, the EA maintains an advantage over COBYLA for both accuracy and variance over an average of 10 runs. In this plot CVaR was used which the reader can note is significantly more stable than using the max_count option. The latter is shown in the bottom plot, where the approximation ratio reaches a value of 1 fewer times, and in general, this method is shown to be less stable. For example, for $n=12$ a particularly hard-to-solve graph appears to have been randomly chosen. Conversely, both optimizers solved graphs with the following two sampled sizes more efficiently. In this plot, the variance for COBYLA is substantially increased and the accuracy takes a large hit, dropping as low as  $\approx\,0.61$ for both $n=20$ and $26$. On the other hand, the EA optimizer is shown to be more robust, even though the accuracy does not remain as high as with CVaR, the variance increases only slightly, and it is still able to maintain an approximation ratio at  $\approx\,0.77$ or above for all cases, with a much lower discrepancy between repetitions than COBYLA. Independently of the fitness function used the EA optimizer always performs on par or better than COBYLA and with a smaller variance
@@
Figure 2: Approximation ratio (3), averaged over 10 runs, for graph sizes $n=[4,10,12,14,16,20,26]$ on $d$-3 regular graphs solved using QAOA paired with either COBYLA or the EA. For sub-figure a) CVaR is used, whilst b) uses max_count to calculate the fitness of solutions. We used $g=10$, $n_{\rm pop}=10$ for all simulations, and the standard deviation for error bars. The horizontal dotted red line represents the maximum approximation ratio achievable.
@@
We theorize this higher robustness for the EA is due to an evolving population of solutions rather than just an individual one, and the fact that the EA is less susceptible to initial conditions, e.g., due to a poor initial guess. Furthermore, the EA can explore a wider search space within the fitness landscape instead of being guided by gradients and is thus able to consistently find more accurate solutions.
@@
Figure 3: Simulated approximation ratio as a function of the population size ($n_{\rm pop}$) for 20-node (left) and 26-node (right) graphs. Simulations were run for values of $n_{\rm pop}=[8,10,12,16,20]$, indicated by the vertical dotted lines with labels. For a given $n_{\rm pop}$, the approximation ratio for single-population runs (black) is compared to multi-population-EA results for $g_{\rm f}=5$ (red) and $g_{\rm f}=7$ (green). Markers represent the mean approximation ratio from 10 independent regular $d$-3 random graphs and the error bars span the minimum and maximum obtained for each $n_{\rm pop}$. The black and green markers are slightly shifted along the $x$-axis to facilitate the comparison; as guidelines, we show solid horizontal lines (and grey boxes) next to the black markers to compare the mean multi-population ratios (and the extent) with the single-population results. The hatched area for ratios above 1 is by definition excluded.
@@
In Fig. 3 we show the simulations for finding the average Max-Cut approximation ratio on a graph with $n=20$ nodes on the left, and $n=26$ on the right, for a single population algorithm versus a multi-population one as explained in the methodology. Note that the population size is the same for each isolated population, thus the MGA approach is effectively evolving double the number of individuals independently. In the left plot, the multi-population algorithm for $g_{\rm f}=5$, shown in red, achieves a slightly better solution quality than the single population for all combinations of population size, over an average of 10 random graphs. For both the single and multi-strategy the mean approximation ratio tends to increase as the population size increases, except for $n_{\rm pop}=16$. This is expected as evaluating more individuals and increasing the population size, should help, but is not necessarily a linear relationship and is considered a tunable parameter. For $g_{\rm f}=7$ this advantage slightly increases, over the one achieved with $g_{\rm f}=5$, in four out of the 5 instances. Beyond achieving a better mean solution, the multi-population strategies also achieve a higher minimum for all experiments. Another interesting thing to observe is that the multi-population algorithm, for both $g_{\rm f}$ values, will in most cases shown, achieve at least the same accuracy rate as the single-population algorithm with a larger population size. This is except for $n_{\rm pop}=16$ again, yet, in the case of $n_{\rm pop}=12$ the multi-population approach with $g_{\rm f}=5$ achieves an accuracy better than the single population one with $n_{\rm pop}=20$. This result suggests there is potential for a speed-up for the MGA approach, but such an investigation of that performance is outside the scope of this paper.
@@
On the right plot instead, the results and advantages are not as simple to interpret. In this case, for $n=26$, there needs to be some tuning between $n_{\rm pop}$ and $g_{\rm f}$ to achieve the best approximation ratio. For the multi-population with $g_{\rm f}=5$, we still achieve better or on-par results, with a lower population size than the single population does with $n_{\rm pop}=20$. A larger $g_{\rm f}$ value only helps with larger population sizes, otherwise, it is detrimental. In these cases, for $n_{\rm pop}\leq 12$, waiting for a larger number of generations might hinder successful evolution and convergence as we are introducing a new fresh individual when the algorithm has already started converging towards another set of genes. This competition between two very fit individuals might split a low-size population in a way that does not explore the fitness landscape optimally. In both plots, the multi-population algorithm with $g_{\rm f}=7$ and $n_{\rm pop}=20$ achieves the highest average approximation ratio overall.
@@ IV-B Hardware
Following the simulation results, we performed an initial experiment on the ibm_nazca device, solving the Max-Cut problem on 5 random $d$-3 regular graphs of size $n=16$. Consistent with simulation, the optimal Max-Cut was found in all cases. Given this, we scaled up the experiments to graphs of size $n=20$ with a population size of $n_{\rm pop}=12$. This test case was chosen because it gave some of the most promising simulation results for the multi-population strategy. As can be seen in Fig. 4, we tested the single and multi-population approach with $g_{\rm f}=5$. Due to constraints on QPU time, we ran these on the hardware for 15 generations rather than 20, and tested on 5 random graphs rather than 10. While this limits any claims we can make regarding the advantages of E-QAOA and multi-population strategies, we present the hardware results nonetheless because they serve as a proof-of-concept and validation of our simulations. The multi-population strategy achieves a slightly higher mean and a higher maximum on hardware than the single population. Both methods perform slightly worse on hardware than in simulation. However, this is expected and it is promising to see that the gap between hardware and simulation is relatively small, and that the small improvement of MGA over single population is maintained on hardware.
@@
Figure 4: As in Fig. 3, but for hardware runs (empty markers) of 5 random regular $d$-3 graphs with $n=20$. Hardware runs configured with $n_{\rm pop}=12$, $g=15$ and $g_{\rm f}=5$. The single-population run (empty square) was performed on ibm_nazca and the multi-population run (empty triangle) was performed on both ibm_nazca and ibm_cusco. For comparison, we show simulation results for $n_{\rm pop}=12$ and 20 as filled markers.
@@
Figure 5: Both figures show a uniqueness ratio (10) over the evolution of generations for a single-population (solid red) and a multi-population (dashed blue) on the average of 5 random graphs of size $n=20$. For all hardware runs $g=15$, $g_{\rm f}=5$ and $n_{\rm pop}=12$. These statistics are achieved on the ibm_nazca machine, with the shaded area representing the standard deviation. Sub-figure a) shows the uniqueness in fitness values calculated with CVaR. Sub-figure b) shows the uniqueness in the first allele of the genetic code, i.e. $\beta_{0}$. Note that for the multi-population approach, the uniqueness is calculated only on one population, within the ibm_nazca machine, as such the values between the two approaches overlap before the first migration. The vertical dotted lines represent the generation number at which these migrations happen. Finally, we use a spline function to smooth the uniqueness ratio values between (discrete) generations.
@@
As mentioned in the introduction a multi-population approach serves to amplify the diversity within the genes of a population, avoid premature convergence, and search the fitness landscape for values that might not otherwise be reached with a single population. Fig. 5 shows how the uniqueness for two different parameters evolves for the single and multi-population approaches as tested on hardware. On the left side, sub-figure (a) shows the uniqueness with respect to how many fitness values are found within the ibm_nazca population. In this instance, the multi-population approach trends slightly higher whilst inter-crossing with the single-population approach. Given that the fitness value is not related in a one-to-one fashion to the genes or ansatz parameters of the circuit, the fitness value cannot show the whole picture on its own, since a family of $\beta$ and $\gamma$ combinations would produce the same CVaR output when evaluating the fitness. To provide more complete information, we also show the uniqueness of $\beta_{0}$ values in the right sub-figure (b). In this case, the separation is more pronounced, with the migrations, shown in the vertical dotted lines happening at regular intervals, pushing and maintaining a higher uniqueness ratio within the multi-population approach. The comparison is done using only one allele as the cross-over operator for recombination, which is happening on all alleles, should generate unique individuals in these terms without the aid of mutation. As for the uniqueness of fitness values, diversity between genes does not provide the whole picture either, and would require further analysis of the expressivity of the circuit for the genes, and a fitness value using only the best Max-Cut value within the distribution, i.e. $\alpha=10^{-4}$, or 1 count in $10^{4}$ shots.
@@
Given these results, we can conclude that the multi-population E-QAOA has potential but may need careful parameter tuning to produce optimal results and successfully compete with the single population E-QAOA strategy. In this respect, we would allow further generations to evolve and converge without migrations once successful swaps have already modified the distributed populations. This further step is necessary to ultimately allow diversity within the population to fall, and let individual solutions converge towards the best genetic codes.
@@ V Conclusion
In this paper, we presented results comparing an evolutionary approach against a standard QAOA using COBYLA, on a simulator for $d$-3 regular graphs, ranging in size from 4 to 26 nodes. We studied the accuracy and variance for the approximation ratio, using both CVaR and max_count, accounting for the former being a more stable fitness evaluator. Next, we noted the best EA parameters to take forward in producing a multi-population E-QAOA, testing this MGA approach both on the simulator and hardware. On the simulator, we tested two different migration schedules to see how these would affect performance and then tested the multi-population E-QAOA with $g_{\rm f}=5$ on the hardware. Results between simulators and hardware were in accordance, even though, as expected, hardware results produced poorer accuracy rates. Lastly, we investigated the uniqueness of fitness and genetic code within a population using a single- and multi-population approach to guarantee an increase in genetic diversity in the latter strategy.
Our experiments, although performed on small-size problems with limited computational resources, have paved the way for future work. Despite these limitations, our work presents a novel approach by using independent isolated populations to evolve E-QAOA and distributing them on 2 scheduled QPUs by running experiments on quantum hardware. Further investigations of multi-population E-QAOA approaches are already underway. These upcoming results aim to consolidate the novel techniques presented in this paper, characterize their advantages, and further probe parallelization techniques. We have set the foundations for this work to eventually produce advancements in speed-ups alongside also boosting accuracy.

### v3

@@ III Set-up
The simulations are conducted on the Scafell Pike high-performance computing system at Sci-Tech Daresbury (UK), which features dual-socket compute units with 16-core Intel Xeon Gold CPUs per socket. Hardware experiments are executed on IBM’s 127-qubit Eagle processors ibm_cusco and ibm_nazca, which exhibit an Error-Per-Layered-Gate (EPLG) on a 100-qubit chain of 5.9% and 3.2%, respectively at the time when the calculations are executed [41]. For each fitness evaluation, the QAOA circuit is sampled with $10^{4}$ shots. We tested experiments (not shown here) varying the number of shots and found that $10^{4}$ provides a suitable convergence for our problem size with the quantum hardware available. For reproducibility, all experimental parameters are summarized in Table I and remain unchanged across experiments.
TABLE I: Summary of variables used in our E-QAOA implementation, their description, and values.
@@
$\gamma$, $\beta$ Ansatz parameters to optimize $[-\pi$, $\pi]$
@@ IV Results and Discussion
Now, we present the results of our experiments and a discussion on their performance. Tests are first run on a classical state vector simulator to evaluate potential advantages and tune the appropriate parameters, ahead of the hardware experiments. Then, they are run on ibm_cusco and ibm_nazca, the IBM Quantum devices available to the authors via the cloud.
@@ IV-A Simulations
Fig. 2 shows the performance of our EA optimizer compared to COBYLA’s for $d$-3 regular graphs sized between 4 and 26 nodes. To ensure fair comparison between the methods, we compute the average and standard deviation of 10 independent COBYLA-based runs with E-QAOA runs with a population size $n_{pop}=10$. In the top panel, we note that both methods achieve the optimal solution, i.e. an approximation ratio of 1 (dotted horizontal line), for small graphs, as expected. However, from $n\geq 16$, both optimizers show a decline in accuracy: the EA maintains an advantage over COBYLA for accuracy and variance. Fig. 2 clearly shows that the CVaR method is more stable than max_count, as expected from prior considerations. The results with max_count are shown in the bottom panel, where the approximation ratio reaches 1 fewer times than for the CVaR case. The $n=12$ graph is an interesting case, where the max_count only achieves an approximation ratio of $\approx 0.5$, compatible with random guessing. The variance for COBYLA is substantially larger than for E-QAOA, and the mean accuracy appears lower than our method (  $\approx\,0.61$ for both $n=20$ and $26$). Conversely, the EA optimizer appears more robust, even though the accuracy does not remain as high as with CVaR, the variance increases only slightly, and it is still able to maintain an approximation ratio at  $\approx\,0.77$ or above for all cases, with a much lower discrepancy between repetitions than COBYLA. Note that the function max_count is equivalent to using the CVaR function with $\alpha=1/\text{shots}$, such that reducing $\alpha$ drastically will create instability in convergence. Through experimentation we found the value of $\alpha=0.15$ to be robust enough for our experiments.
@@
Figure 2: Approximation ratio (Eq. 3) averaged over 10 runs per graph, for graph sizes $n=[4,10,12,14,16,20,26]$ on $d$-3 regular graphs solved using QAOA paired with either COBYLA (green) or the EA (purple). We show results using CVaR (top) and max_count (bottom) for the fitness of solutions. In the EA, we set $g=10$ and $n_{\rm pop}=10$ and the COBYLA-based runs are computed from 10 independent runs run for 10 iterations. The error bars indicate the standard deviation of 10 realizations. The horizontal dotted line represents the maximum approximation ratio achievable.
@@
We expect that the higher robustness for the EA is due to an evolving population of solutions rather than just an individual realization; additionally, our runs appear to suggest that the EA is less susceptible to initial conditions, e.g., due to a poor initial guess. Furthermore, the EA can explore a wider search space within the fitness landscape instead of being guided by gradients and is thus able to consistently find more accurate solutions.
@@
Figure 3: Simulated approximation ratio as a function of the population size ($n_{\rm pop}$) for 20-node (left) and 26-node (right) graphs. Classical simulations were run for values of $n_{\rm pop}=[8,10,12,16,20]$, indicated by the vertical dotted lines with labels. For a given $n_{\rm pop}$, the approximation ratio for single-population runs (black) is compared to multi-population-EA results for $g_{\rm f}=5$ (red) and $g_{\rm f}=7$ (green). Markers represent the average approximation ratio from 10 distinct independent regular $d$-3 random graphs, and the error bars span the minimum and maximum obtained for each $n_{\rm pop}$. The black and green markers are slightly shifted along the $x$-axis to facilitate the comparison; as guidelines, we show solid horizontal lines (and grey boxes) next to the black markers to compare the mean multi-population ratios (and the extent) with the single-population results. The hatched area for ratios above 1 is excluded.
@@
In Fig. 3, we show the simulations for finding the average Max-Cut approximation ratio on a graph with $n=20$ nodes on the left, and $n=26$ on the right panel for a single-population algorithm versus a multi-population one as explained in the methodology. Note that the population size is the same for each isolated population; therefore, the MGA approach is effectively evolving double the number of individuals independently. In the left panel, the multi-population algorithm for $g_{\rm f}=5$, shown in red, achieves a slightly better solution quality than the single population for all combinations of population size, over an average of 10 random graphs. For both the single and multi-strategy, the mean approximation ratio tends to increase as the population size increases, except for $n_{\rm pop}=16$. This is expected as evaluating more individuals by increasing the population size should enhance convergence, albeit not necessarily in a linear fashion. For $g_{\rm f}=7$, this advantage slightly increases over the one achieved with $g_{\rm f}=5$, in four out of the 5 instances. Beyond achieving a better mean solution, the multi-population strategies also achieve a higher minimum for all experiments. Another interesting thing to observe is that the multi-population algorithm, for both $g_{\rm f}$ values, will, in most cases shown, achieve at least the same accuracy rate as the single-population algorithm with a larger population size. This is except for $n_{\rm pop}=16$ again, yet, in the case of $n_{\rm pop}=12$ the multi-population approach with $g_{\rm f}=5$ achieves an accuracy better than the single population one with $n_{\rm pop}=20$. This result suggests there is potential for a speed-up for the MGA approach, but such an investigation of that performance is outside the scope of this paper.
@@
On the other hand, the right panel illustrating results for $n=26$ suggests that further tuning is required between $n_{\rm pop}$ and $g_{\rm f}$ to achieve better approximation ratios. For the multi-population with $g_{\rm f}=5$, we still achieve better or on-par results, with a lower population size than the single population does with $n_{\rm pop}=20$. A larger $g_{\rm f}$ value only helps with larger population sizes, otherwise, it is detrimental. In these cases, for $n_{\rm pop}\leq 12$, waiting for a larger number of generations might hinder successful evolution and convergence as we introduce a new, fresh individual when the algorithm has already started converging towards another set of genes. This competition between two very fit individuals might split a small population in a way that does not explore the fitness landscape optimally. In both plots, the multi-population algorithm with $g_{\rm f}=7$ and $n_{\rm pop}=20$ achieves the highest average approximation ratio overall.
@@ IV-B Hardware
Following promising simulation results, we validate our approach experimentally on the ibm_nazca quantum device by solving the Max-Cut problem for 5 random $d$-3 regular graphs of size $n=16$. As in the simulation, the optimal Max-Cut solution is found in all cases. Subsequently, we scale to graphs with $n=20$ vertices and a population size of $n_{\rm pop}=12$. This system is selected because it produced the highest approximation ratio and quickest convergence for classical state vector simulations multi-population strategy. We show results for a single and multi-population approach with $g_{\rm f}=5$ in Fig. 4. The multi-population strategy achieves a slightly higher mean and a higher maximum on hardware than the single population. Both methods perform slightly worse on hardware than in simulation. However, this effect is expected, and the small discrepancy between hardware and simulations suggests that (i) the method could be successfully run on larger systems, and (ii) the improvement of MGA over single population is maintained on hardware. Due to constraints on available QPU time, we simulate these systems on the hardware platform for 15 generations rather than 20, and test 5 out of 10 random graphs. While this restriction may mitigate claims regarding the advantages of E-QAOA and multi-population strategies, we present the hardware results as useful proof-of-concept research and a robust validation of our classical simulations.
@@
Figure 4: As in Fig. 3, but for hardware runs (empty markers) of 5 random regular $d$-3 graphs with $n=20$. Hardware runs configured with $n_{\rm pop}=12$, $g=15$ and $g_{\rm f}=5$. The single-population run (empty square) was performed on ibm_nazca and the multi-population run (empty triangle) was performed on both ibm_nazca and ibm_cusco. For comparison, we show simulation results for $n_{\rm pop}=12$ and 20 as filled markers. The average is computed over 5 different graphs.
@@
Figure 5: Both figures show a uniqueness ratio over the evolution of generations for a single-population (solid red) and a multi-population (dashed blue) on the average of 5 random graphs of size $n=20$. For all hardware runs $g=15$, $g_{\rm f}=5$ and $n_{\rm pop}=12$. These statistics are achieved on the ibm_nazca machine, with the shaded area representing the standard deviation. Sub-figure a) shows the uniqueness in fitness values calculated with CVaR. Sub-figure b) shows the uniqueness in the first allele of the genetic code, i.e. $\beta_{0}$. Note that for the multi-population approach, the uniqueness is calculated only on one population, within the ibm_nazca machine; as such, the values between the two approaches overlap before the first migration. The vertical dotted lines represent the generation number at which these migrations happen. Finally, we use a spline function to smooth the uniqueness ratio values between (discrete) generations.
@@
A multi-population approach can enhance the diversity within the genes of a population, avoid premature convergence, and search the fitness landscape for values that might not otherwise be reached with a single population. Also note that the heterogeneity of different devices does not have a negative impact on the evolution of populations, rather it can be used to promote healthy diversity that aids more effective searching of the solution space. Fig. 5 shows how the uniqueness for two different parameters evolves for the single and multi-population approaches as tested on hardware. The left panel (a) shows the uniqueness with respect to how many fitness values are found within the ibm_nazca population. Here, the multi-population approach trends slightly higher whilst intercrossing with the single-population approach. Given that the fitness value is not related in a one-to-one fashion to the genes or ansatz parameters of the circuit, the fitness value cannot show the whole picture on its own, since a family of $\beta$ and $\gamma$ combinations would produce the same CVaR output when evaluating the fitness. To provide more complete information, we also show the uniqueness of $\beta_{0}$ values in the right panel (b). In this case, the separation is more pronounced, with the migrations, shown in the vertical dotted lines happening at regular intervals, pushing and maintaining a higher uniqueness ratio within the multi-population approach. The comparison is performed using only one allele as the cross-over operator for recombination, which occurs on all genes by construction, is expected to generate unique individuals in these terms without mutation. As for the uniqueness of fitness values, neither does the diversity between genes provide the whole picture, requiring further analysis of the expressivity of the circuit for the genes, and a fitness value using only the best Max-Cut value within the distribution, i.e. $\alpha=10^{-4}$, or 1 count in $10^{4}$ shots.
@@
Given these results, we can conclude that the multi-population E-QAOA has potential but may need careful parameter tuning to produce optimal results and successfully compete with the single-population E-QAOA strategy. In this respect, we would allow further generations to evolve and converge without migrations once successful swaps have already modified the distributed populations. This further step is necessary to ultimately allow diversity within the population to fall and let individual solutions converge towards the best genetic codes.
@@ V Conclusion
This work introduces a multi-population evolutionary QAOA framework for the Max-Cut problem, integrating an evolutionary algorithm into QAOA’s variational parameter optimization protocol. In Fig. 2, we benchmark this evolutionary QAOA approach against a standard COBYLA-based QAOA on random regular graphs (up to 26 nodes) in classical simulation, and in Figs. 4 and 3, we run similar calculations on real IBM quantum devices.
The approach consistently attained equal or higher approximation ratios than COBYLA, for a given number of iterations, while yielding significantly lower cost variance, indicating reliable performance and reproducibility. Notably, employing the CVaR (Conditional Value at Risk) metric as the objective expectation from the tail of the sampled distribution enhanced outcome stability and robustness to noise, resulting in more consistent solution accuracy. In Fig. 3, we demonstrate a distributed execution of the multi-population E-QAOA by running two isolated populations in parallel on separate QPUs with periodic migration of elite solutions between them every $g_{\rm f}=5$ generations. This strategy preserved greater genetic diversity and outperformed the single-population baseline (Fig. 5), confirming the viability of parallel population evolution on actual quantum processors. The hardware results align with simulations (aside from expected noise-induced degradation), showing that our method’s advantages carry over to real quantum devices. Overall, our study demonstrates a distributed QAOA paradigm that improves (i) solution accuracy, (ii) reliability over conventional single-optimizer approaches, and (iii) can scale on heterogeneous networks of quantum processors.
These findings lay a foundation for future work on parallel and distributed quantum optimization, such as scaling to larger problem instances beyond 100 qubits on multiple QPUs, refining inter-population communication schemes (see also [42, 43, 44]). Investigations into adaptive evolutionary strategies, such as dynamic migration intervals, topology-aware elite exchanges, and hybrid evolutionary-gradient optimization methods, could further accelerate convergence rates and maintain genetic diversity. Additionally, extending the framework to other combinatorial optimization tasks beyond Max-Cut, such as portfolio optimization, scheduling, and constrained optimization, is a natural next step, given QAOA’s versatility in addressing various QUBO-formulated problems. Finally, results from different quantum hardware technologies and providers should be compared and benchmarked. We envisage significantly more accurate and better-converged results in the upcoming early fault-tolerant regime, where hardware topologies can handle densely connected graphs and sample distributions at a higher frequency ($\approx 200$ kHz for IBM Heron devices).
Finally, proposed distributed E-QAOA is particularly suited for graph-based optimization problems in domains like network design [45], quantum machine learning (feature selection, hyperparameter tuning, or quantum data encoding) [46, 47], quantum chemistry (molecular electronic structure and ground-state energy prediction) [48, 49], and high-energy physics [50, 47]. The combined benefits of parallel search and evolutionary robustness position E-QAOA as a valuable tool in these demanding quantum optimization applications.